A solid plastic sphere of radius 10 cm has charge with uniform density throughout its volume. The electric field 5 cm from the center is 86 kN/C radially inward. Find the magnitude of the electric field 15cm from the center.



1
Expert's answer
2021-05-20T18:22:50-0400

"E_5=\frac{kqr_5}{R^3}\implies q=\frac{E_5 R^3}{kr_5},"

"E_{15}=\frac{kq}{r_{15}^2}=\frac{kE_5R^3}{r_{15}^2kr_5}=\frac{E_5 R^3}{r_{15}^2r_5}=76~\frac{kN}{C}."